# Predictability Enables Parallelization of Nonlinear State Space Models

### v1

@@ C.1 A Generalized Proof that the Largest Lyapunov Exponent Controls the PL Constant
[…]
@@
[…]
@@
Appendix D DEER Merit Function Inherits Lipschitzness of Dynamics
This section provides a proof of main text Theorem˜3.
Theorem (Theorem˜3).
@@ Appendix D DEER Merit Function Inherits Lipschitzness of Dynamics
[…]
@@
[…]
@@ Appendix E DEER Always Converges Linearly
This section provides a proof of Theorem˜4.
@@
[…]
@@
To ensure the right-hand side of (30) does not exceed a prescribed $\beta\in[0,1)$, choose
@@ Appendix F DEER Converges Globally with Small Overshoot for Sufficiently Strongly Contracting Systems
[…]
@@ Appendix G Alternative Descent Techniques & Worst/Average Complexity
Many inexact GN variants achieve global convergence from any starting point. These include step-size schemes that approximate a continuous flow [Gavurin, 1958], trust-regions that bound update size (yielding ELK when applied to DEER [Gonzalez et al., 2024]), and backtracking line search ensuring loss reduction at each step [Nocedal and Wright, 2006].
@@ Appendix H Proof of Size of Basin of Quadratic Convergence
This section provides a proof of Theorem˜5:
Theorem (Theorem˜5).
@@
[…]
@@
[…]
@@ Appendix I Parameterizing Nonlinear SSMs to be contractive
[…]
@@
[…]
@@
Accordingly, when making Figure 2 (Center), we use the monotonic relationship between $g$ and the LLE from Figure 4 to map the average number of DEER steps (over 20 different seeds) needed for convergence for different values of $g$ to the appropriate value of the LLE. We use 50 values of $T$ from 9 to 9999 (log spaced) to make Figure 2 (Center). We then chose the value of $T$ closest to 1000 to highlight in Figure 2 (Right).
@@
i.e. the lower bound on $\mu$ from Theorem˜2, with $a=1$.
@@
First, as discussed in Gonzalez et al. [2024], DEER can experience numerical overflow when deployed on unstable systems. While we reset to the initialization (in this experiment we initialized $\mathbf{s}_{1:T}$ with iid draws from $\mathcal{U}[0,1]$), doing so slows convergence. Thus, many of our runs for $\lambda>0$ and large $T$ take the maximum number of DEER iteration we allow (we do not allow more than $T$ iterations, as this is the theoretical upper bound for number of DEER iterations before convergence, cf. Proposition 1 of [Gonzalez et al., 2024]), which helps to explain the slight increase in red space for experiment (center plot of Figure 2) vs. theory (left plot of Figure 2). Note, however, that for $T=954$ (the sequence length shown in the right plot), there is no numerical overflow for the DEER trajectories for any of the 20 random seeds or 50 values of $g$ tried.
@@
Second, we observe that for many values of $\lambda$ in the chaotic range, even after the maximum number of DEER steps ($T$) was taken, there was still a large discrepancy between the true sequential rollout and the converged DEER iteration, even though the converged DEER iteration had numerically zero merit function. For example, in Figure 2 (Right), there are a series of points in the top right of the graph that all sit on the line $T=954$, and while they have numerically zero merit function value, the converged DEER trajectories are quite different from the true sequential trajectories. The reason for this behavior precisely stems from the fact that for large values of $g$ (equivalently $\lambda$), these mean-field RNNs are chaotic. Even working in FP64, if slight numerical errors are introduced at any time point in the sequence (say $t=1$), then over the sequence length we can observe exponential divergence from the true trajectories, as illustrated in Figure 5. This experimental observation is complemented by our discussion of why unpredictable systems have excessively flat merit functions in Section 3.2, and provides a numerical perspective on why ill-conditioned landscapes are hard to optimize: if the landscape is extremely flat, many potential trajectories $\mathbf{s}_{1:T}$ can have numerically zero merit function, even in extremely high precision.
@@
Figure 5: Chaotic behavior means numerically zero merit function can still be far from sequential trajectory. For $g=1.85$ and $T=954$, we show the final DEER vs sequential trajectory. The DEER trajectory has merit function (2) numerically equal to zero. However: (Left) the mean absolute deviation (MAD) at each time point $t$ between the final DEER iteration $\mathbf{s}_{t}^{(954)}$ and the sequential rollout $\mathbf{s_{t}^{*}}$ grows exponentially. This exponential growth of error is a signature of chaos: compare, for example, with Figure 9.3.5 of Strogatz [2018]. The saturation of the error eventually occurs because of the saturating nonlinearity present in the RNN. (Right) We visualize the first coordinate of both the final DEER iteration and the sequential trajectory, showing that while they initially coincide, they diverge around $t=300$.
@@
In this section, we provide further experiments in the setting of the mean-field RNN (Figure˜2). In particular, we showcase the generality of our theory beyond DEER (Gauss-Newton optimization), and the practicality of our theory by reporting wallclock times. We consider the setting in the right most panel of Figure˜2, where we evaluate a mean field RNN over a sequence length of length $T=954$.
@@
We include a quasi-Newton algorithm proposed in Gonzalez et al. [2024] called quasi-DEER. Quasi-DEER simply replaces the $J_{t}$ defined in eq.˜4 with $\mathrm{diag}(J_{t})$, and so is also parallelizable over the sequence length with a parallel scan. Furthermore, we also include gradient descent on the merit function, which is embarrassingly parallel over the sequence length. In the top panel of Figure˜6, we observe that the number of steps for gradient descent and quasi-DEER to converge also scales monotonically with the LLE, as we expect from Theorem˜2. DEER (Gauss-Newton) converges in a small number of steps all the way up to the threshold between predictability and unpredictability ($\lambda=0$). Intuitively, the performance of the other optimizers degrades more quickly as unpredictability increases because quasi-Newton and gradient descent use less information about the curvature of the loss landscape.
Even though gradient descent was slower to converge in this setting, we only tried gradient descent with a fixed step size. An advantage of a first-order method like gradient descent over a second-order method like Gauss-Newton (DEER) is that the first-order method is embarrassingly parallel (and so with sufficient parallel processors, the update runs in constant time), while DEER and quasi-DEER use parallel scans (and so the update runs in $O(\log T)$ time). Exploring accelerated first-order methods like Adam [Kingma and Ba, 2015], or particularly Shampoo [Gupta et al., 2018] or SOAP [Vyas et al., 2025] (which are often preferred in recurrent settings like eq.˜1)—or in general trying to remove the parallel scan—are therefore very interesting directions for future work.
[…]
@@
We run the experiment in Figure˜6 on a smaller scale than the experiment in Figure˜2 (Right). In Figure˜6, we consider 5 random seeds for 16 values of $g$ equispaced between $0.5$ and $2.0$. Each wallclock time reported is the average of 5 runs for the same seed. While DEER (Gauss-Newton) and quasi-DEER effectively do not have a step size (they use a step size of $1$ always). For each value of $g$, we ran gradient descent with the following set of step sizes $\alpha$: $0.01,0.1,0.25,0.5,0.6,0.7,0.8,0.9,\text{and }1.0$. For each value of $g$, we then pick the step size $\alpha$ that results in the fastest convergence of gradient descent. For the smallest value of $g=0.5$, we use $\alpha=0.6$; for $g=0.6$, we use $\alpha=0.5$; and for all other values of $g$, we use $\alpha=0.25$. Future work may investigate more adaptive ways to tune the step size $\alpha$, or to use a learning rate schedule.
We use a larger tolerance of $\mathcal{L}(\mathbf{s})\leq 0.1$ to declare convergence than in the rest of the paper (where we use a tolerance of $10^{-7}$) because gradient descent often did not converge to the same degree of numerical precision as sequential, quasi-DEER, or DEER. However, a squared error of $0.1$ over a sequence length of length $T=954$ is equivalent to a per time-step average error on the order of $10^{-4}$, in a system where $D=100$ and each state has current on the order of $1$. Nonetheless, it is an interesting direction for future work to investigate how to get gradient descent to converge to greater degrees of numerical precision in these settings; and, in general, how to improve the performance of all of these parallel sequence evaluators in lower numerical precision.
@@
We form the two-well potential for our experiment in Section 5 as a sum of two quadratic potentials. Concretely, we define the potential $\phi$ as the negative log probability of the mixture of two Gaussians, where one is centered at (0,1.4)(0,-1.4) and the other is centered at $(0,1.6)$, and they both have diagonal covariance. In Langevin dynamics [Langevin, 1997, Friedman, 2022] for a potential $\phi$, the state $s_{t}$ evolves according to
@@
where $\epsilon$ is the step size and $w_{t}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}\mathcal{N}(0,I_{D})$. Accordingly, the Jacobians of the dynamics (those used in DEER) take the form
@@
Moreover, a further added benefit of our theory is demonstrated by our choice of initialization of DEER. Both [Lim et al., 2024] and [Gonzalez et al., 2024] exclusively initialized all entries of $\mathbf{s}^{(0)}$ to zero. However, such an initialization can be extremely pathological if the region of state space containing $\mathbf{0}$ is unstable, as is the case for the particular two well potential we consider. For this reason, we initialize $\mathbf{s}^{(0)}$ at random (as iid standard normals).
An important consequence of this experiment is that it shows that there are systems that are not globally contracting that nonetheless enjoy fast rates of convergence with DEER. This fact is important because a globally contractive neural network may not be so interesting/useful for classification, while a locally contracting network is.
Figure 7: In this plot, we provide additional information about the behavior of DEER when rolling out Langevin dynamics on a two-well potential. (Left) We observe that across 20 random seeds (including different Langevin dynamics trajectories), the LLE for intermediate DEER iterations becomes negative after the first iteration. Consequently, we observe that the merit function (Center) experiences a spike on the very first DEER iteration (following initialization, which was the only trajectory with positive LLE), before trending towards convergence. As the system spends most of its time in contracting regions, we observe (Right) that the number of DEER iterations needed for convergence scales sublinearly with the sequence length $T$. We plot the min-max range for 20 seeds, and observe that even out of 20 seeds, the maximum number of DEER iterations needed to converge on a sequence length of $T=10,000$ is just more than $30$.
@@
To further demonstrate the applicability of our results—and to validate them in the context of non-autonomous systems—we construct nonlinear observers. Observers are commonly used in science and engineering to reconstruct the full state of a system from partial measurements [Luenberger, 1979, Simon, 2006]. As a benchmark, we consider nine chaotic flows from the dysts dataset [Gilpin, 2021b]. According to Theorem (2), these systems exhibit poorly conditioned merit function landscapes and are thus not well-suited for parallelization via DEER. If the corresponding observers are stable, then they should be suitable for DEER.
@@
J.6 Numerical computation of the LLE

### v2

@@ C.1 A Generalized Proof that the Largest Lyapunov Exponent Controls the PL Constant
As discussed in the main text, the predictability of the nonlinear state space model is characterized by the products of its Jacobians along a trajectory. We will need to control how this product behaves. To reduce notational burden, we will drop the DEER iteration superscript $i$. In particular, we will assume that there exists a function $g_{J}:\mathbb{N}_{0}\to\mathbb{R}$ such that
@@
Let $\|\cdot\|_{\xi}$ be the matrix operator norm induced by the vector norm $\|\cdot\|_{\xi}$. Suppose there is a function $g_{J}:\mathbb{N}_{0}\to\mathbb{R}$ such that
@@
Appendix D DEER Merit Function Inherits Lipschitzness of Dynamics Jacobians
This section provides a proof of main text ˜3.
Theorem (˜3).
@@ Appendix D DEER Merit Function Inherits Lipschitzness of Dynamics Jacobians
$\forall\,t>1,\ \ \ s,s^{\prime}\in\mathbb{R}^{D}:\quad\|J_{t}(s)-J_{t}(s^{\prime})\|\leq L\|s-s^{\prime}\|,$
@@
$\forall s,s^{\prime}\in\mathbb{R}^{D}:\quad\|J_{t}(s_{t})-J_{t}(s_{t}^{\prime})\|_{2}\;\leq\;L\,\|s_{t}-s_{t}^{\prime}\|_{2}.$
@@ Appendix E DEER Always Converges Linearly
This section provides a proof of ˜4.
@@
$\mathbf{W}\;\coloneqq\;\operatorname{Diag}\bigl(I_{D},\;w^{2}I_{D},\;\dots,\;w^{2T}I_{D}\bigr)\;\in\;\mathbb{R}^{TD\times TD},\qquad w>0.$ (27)
@@
To ensure the right‑hand side of (30) does not exceed a prescribed $\beta\in[0,1)$, choose
@@ Appendix F DEER Converges Globally with Small Overshoot for Sufficiently Strongly Contracting Systems
In this section we show that DEER converges globally to the optimum $\mathbf{s}^{*}$ when the nonlinear state space model (1) is sufficiently strongly contracting. To do so, we first briefly recall the assumptions of Lemma 1. Let $\|\cdot\|_{\xi}$ be the matrix operator norm induced by the vector norm $\|\cdot\|_{\xi}$. Suppose there is a function $g_{J}:\mathbb{N}_{0}\to\mathbb{R}$ such that
@@ Appendix G Alternative Descent Techniques & Worst/Average Complexity
Many inexact GN variants achieve global convergence from any starting point. These include step-size schemes that approximate a continuous flow [76], trust-regions that bound update size (yielding ELK when applied to DEER [3]), and backtracking line search ensuring loss reduction at each step [19].
@@ Appendix H Proof of Size of Basin of Quadratic Convergence
This section provides a proof of ˜5:
Theorem (˜5).
@@
Suppose we are at a point $\mathbf{s}^{(i)}\in\mathbb{R}^{TD}$ (i.e. DEER iterate $i$), and we want to get to $\mathbf{s}^{(i+1)}$. The change in the trajectory obtained from eq.˜3 is,
@@
which is the direct analogy of Boyd and Vandenberghe [28, 9.33]. To reiterate, here $L$ is the Lipschitz constant of $\mathbf{J}$, while $\mu:=\inf_{i\in\mathbb{N}}\sigma^{2}_{\mathrm{min}}\left(\mathbf{J}(\mathbf{s}^{(i)})\right)$.
@@ Appendix I Parameterizing Nonlinear SSMs to be contractive
All of our experiments use FP64 to, as much as possible, focus on algorithmic factors controlling the rate of convergence of DEER, as opposed to numerical factors. As noted in [3], DEER can be prone to numerical overflow in lower precision. While such numerical overflow can be overcome by resetting NaNs to their initialized value, such an approach resets the optimization and leads to rates that are slower than what Gauss-Newton would achieve in infinite precision (exact values in $\mathbb{R}$).
@@
for mild sinusoidal inputs $u_{t}$. We have $s_{t}\in\mathbb{R}^{D}$, where in our experiments $D=100$. Note that because of the placement of the saturating nonlinearity, here $s_{t}$ represents current, not voltage.
@@
Accordingly, when making Figure 2 (Center), we use the monotonic relationship between $g$ and the LLE from Figure 4 to map the average number of DEER steps (over 20 different seeds) needed for convergence for different values of $g$ to the appropriate value of the LLE. We use 50 values of $T$ from 9 to 9999 (log spaced) to make Figure 2 (Center). We highlight $T=1000$ in Figure 2 (Right).
@@
i.e. the lower bound on $\mu$ from ˜2, with $a=1$.
@@
First, as discussed in Gonzalez et al. [3], DEER can experience numerical overflow when deployed on unstable systems. While we reset to the initialization (in this experiment we initialized $\mathbf{s}_{1:T}$ with iid draws from $\mathcal{U}[0,1]$), doing so slows convergence. Thus, many of our runs for $\lambda>0$ and large $T$ take the maximum number of DEER iteration we allow (we do not allow more than $T$ iterations, as this is the theoretical upper bound for number of DEER iterations before convergence, cf. Proposition 1 of [3]), which helps to explain the slight increase in red space for experiment (center plot of Figure 2) vs. theory (left plot of Figure 2). Note, however, that for $T=1000$ (the sequence length shown in the right plot), there is no numerical overflow for the DEER trajectories for any of the 20 random seeds or 50 values of $g$ tried.
@@
Second, we observe that for many values of $\lambda$ in the chaotic range, even after the maximum number of DEER steps ($T$) was taken, there was still a large discrepancy between the true sequential rollout and the converged DEER iteration, even though the converged DEER iteration had numerically zero merit function. For example, in Figure 2 (Right), there are a series of points in the top right of the graph that all sit on the line $T=1000$, and while they have numerically zero merit function value, the converged DEER trajectories are quite different from the true sequential trajectories. The reason for this behavior precisely stems from the fact that for large values of $g$ (equivalently $\lambda$), these mean-field RNNs are chaotic. Even working in FP64, if slight numerical errors are introduced at any time point in the sequence (say $t=1$), then over the sequence length we can observe exponential divergence from the true trajectories, as illustrated in Figure 5. This experimental observation is complemented by our discussion of why unpredictable systems have excessively flat merit functions in Section 3.2, and provides a numerical perspective on why ill-conditioned landscapes are hard to optimize: if the landscape is extremely flat, many potential trajectories $\mathbf{s}_{1:T}$ can have numerically zero merit function, even in extremely high precision.
@@
Figure 5: Chaotic behavior means numerically zero merit function can still be far from sequential trajectory. For $g=1.85$ and $T=1000$, we show the final DEER vs sequential trajectory. The DEER trajectory has merit function (2) numerically equal to zero. However: (Left) the mean absolute deviation (MAD) at each time point $t$ between the final DEER iteration $\mathbf{s}_{t}^{(T)}$ and the sequential rollout $\mathbf{s_{t}^{*}}$ grows exponentially. This exponential growth of error is a signature of chaos: compare, for example, with Figure 9.3.5 of Strogatz [10]. The saturation of the error eventually occurs because of the saturating nonlinearity present in the RNN. (Right) We visualize the first coordinate of both the final DEER iteration and the sequential trajectory, showing that while they initially coincide, they diverge around $t=300$.
@@
In this section, we provide further experiments in the setting of the mean-field RNN (Figure˜2). In particular, we showcase the generality of our theory beyond DEER (Gauss-Newton optimization), and the practicality of our theory by reporting wallclock times. We consider the setting in the right most panel of Figure˜2, where we evaluate a mean field RNN over a sequence length of length $T=1000$.
@@
We include a quasi-Newton algorithm proposed in Gonzalez et al. [3] called quasi-DEER. Quasi-DEER simply replaces the $J_{t}$ defined in eq.˜4 with $\mathrm{diag}(J_{t})$, and so is also parallelizable over the sequence length with a parallel scan. Furthermore, we also include gradient descent on the merit function, which is embarrassingly parallel over the sequence length. In the top panel of Figure˜6, we observe that the number of steps for gradient descent and quasi-DEER to converge also scales monotonically with the LLE, as we expect from ˜2. DEER (Gauss-Newton) converges in a small number of steps all the way up to the threshold between predictability and unpredictability ($\lambda=0$). Intuitively, the performance of the other optimizers degrades more quickly as unpredictability increases because quasi-Newton and gradient descent use less information about the curvature of the loss landscape.
Even though gradient descent was slower to converge in this setting, we only tried gradient descent with a fixed step size. An advantage of a first-order method like gradient descent over a second-order method like Gauss-Newton (DEER) is that the first-order method is embarrassingly parallel (and so with sufficient parallel processors, the update runs in constant time), while DEER and quasi-DEER use parallel scans (and so the update runs in $O(\log T)$ time). Exploring accelerated first-order methods like Adam [77], or particularly Shampoo [78] or SOAP [79] (which are often preferred in recurrent settings like eq.˜1)—or in general trying to remove the parallel scan—are therefore very interesting directions for future work.
We note that sequential evaluation of eq.˜1 can also be thought of as block coordinate descent on the merit function $\mathcal{L}(\mathbf{s})$, where the block $s_{t}\in\mathbb{R}^{D}$ is optimized at optimization step $(t)$. The optimization of each block is a convex problem: simply minimize $\|s_{t}-f(s_{t-1}^{*})\|_{2}^{2}$, or equivalently set $s_{t}=f(s_{t-1}^{*})$. As sequential evaluation will always take $T$ steps to converge, we do not include it in the top panel of Figure˜6.
@@
We run the experiment in Figure˜6 on a smaller scale than the experiment in Figure˜2 (Right). In Figure˜6, we consider 5 random seeds for 16 values of $g$ equispaced between $0.5$ and $2.0$. Each wallclock time reported is the average of 5 runs for the same seed. We use a batch size of 1. While DEER (Gauss-Newton) and quasi-DEER effectively do not have a step size (they use a step size of $1$ always). For each value of $g$, we ran gradient descent with the following set of step sizes $\alpha$: $0.01,0.1,0.25,0.5,0.6,0.7,0.8,0.9,\text{and }1.0$. For each value of $g$, we then pick the step size $\alpha$ that results in the fastest convergence of gradient descent. For the smallest value of $g=0.5$, we use $\alpha=0.6$; for $g=0.6$, we use $\alpha=0.5$; and for all other values of $g$, we use $\alpha=0.25$. Future work may investigate more adaptive ways to tune the step size $\alpha$, or to use a learning rate schedule.
We use a larger tolerance of $\nicefrac{{\mathcal{L}(\mathbf{s})}}{{T}}\leq 10^{-4}$ to declare convergence than in the rest of the paper (where we use a tolerance of $10^{-10}$) because gradient descent often did not converge to the same degree of numerical precision as sequential, quasi-DEER, or DEER. However, this is a per time-step average error on the order of $10^{-4}$, in a system where $D=100$ and each state has current on the order of $1$. Nonetheless, it is an interesting direction for future work to investigate how to get gradient descent to converge to greater degrees of numerical precision in these settings; and, in general, how to improve the performance of all of these parallel sequence evaluators in lower numerical precision.
@@
We form the two-well potential for our experiment in Section 5 as a sum of two quadratic potentials. Concretely, we define the potential $\phi$ as the negative log probability of the mixture of two Gaussians, where one is centered at (0,1.4)(0,-1.4) and the other is centered at $(0,1.6)$, and they both have diagonal covariance. In Langevin dynamics [80, 81] for a potential $\phi$, the state $s_{t}$ evolves according to
@@
where $\epsilon$ is the step size and $w_{t}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}\mathcal{N}(0,I_{D})$.  [^3] Notice that this is a discretization (with time step $\epsilon$) of the Langevin Diffusion SDE $ds(t)=-\nabla\phi(s(t))dt+\sqrt{2}dw(t)$, where $w(t)$ is Brownian motion [82, 83, 84]. Accordingly, the Jacobians of the dynamics (those used in DEER) take the form
@@
Moreover, a further added benefit of our theory is demonstrated by our choice of initialization of DEER. Both [1] and [3] exclusively initialized all entries of $\mathbf{s}^{(0)}$ to zero. However, such an initialization can be extremely pathological if the region of state space containing $\mathbf{0}$ is unstable, as is the case for the particular two well potential we consider. For this reason, we initialize $\mathbf{s}^{(0)}$ at random (as iid standard normals).
An important consequence of this experiment is that it shows that there are systems that are not globally contracting that nonetheless enjoy fast rates of convergence with DEER. This fact is important because a globally contractive neural network may not be so interesting/useful for classification, while a locally contracting network could be.
Figure 7: In this plot, we provide additional information about the behavior of DEER when rolling out Langevin dynamics on a two-well potential. (Left) We observe that across 20 random seeds (including different Langevin dynamics trajectories), the LLE for intermediate DEER iterations becomes negative after the first iteration. Consequently, we observe that the merit function (Center) experiences a spike on the very first DEER iteration (following initialization, which was the only trajectory with positive LLE), before trending towards convergence. As the system spends most of its time in contracting regions, we observe (Right) that the number of DEER iterations needed for convergence scales sublinearly with the sequence length $T$. We plot the min-max range for 20 seeds, and observe that even out of 20 seeds, the maximum number of DEER iterations needed to converge on a sequence length of $T=10,000$ is around $35$.
@@
To further demonstrate the applicability of our results—and to validate them in the context of non-autonomous systems—we construct nonlinear observers. Observers are commonly used in science and engineering to reconstruct the full state of a system from partial measurements [30, 31]. As a benchmark, we consider nine chaotic flows from the dysts dataset [85]. According to Theorem (2), these systems exhibit poorly conditioned merit function landscapes and are thus not well-suited for parallelization via DEER. If the corresponding observers are stable, then they should be suitable for DEER.
@@
J.6 Numerical computation of the discrete-time LLE